OpenStudy (anonymous):

Please help?! A projectile is thrown upward so that its distance above the ground after t seconds is given by the function h(t) = - 16t2 + 704t. After how many seconds does the projectile take to reach its maximum height? Show your work for full credit.

OpenStudy (campbell_st):

several ways to do the question.. 1. using calculus 2. using algebra, which i"ll use the max height is on the line of symmetry and the time that is the line of symmetry is found using \[t = \frac{-b}{2a}\] in your question a = -16 and b = 704 substitute then evaluate hope this helps

OpenStudy (campbell_st):

you can find the value of the max height by substituting the time vale from above into the equation.

OpenStudy (anonymous):

Oh, thanks. You're awesome. c:

OpenStudy (anonymous):

So it would be -22? @campbell_st And anytime it says something like that, bout maximun height, it's asking about the vertex? I feel dumb.

OpenStudy (campbell_st):

no its \[t = \frac{-704}{2 \times -16} = \frac{704}{32}\] the negative in the numerator divided by the negative in the denominator make it a positive... so it 22 seconds

OpenStudy (campbell_st):

so the max height occurs at t =22 you can substitute t = 22 to find the max height..., which will be the vertex the vertex is in the line of symmetry.... t = 22 so the vertex is (22, h(22)

OpenStudy (anonymous):

Yes, sorry. I did put positive 22. Thanks you guys! God bless! c: